A 3kg object falls off an 9 m tall ledge.Due to friction it is only going 10 m/s at the ground.Calculate efficiency

g=9.8
1
Expert's answer
2020-10-30T13:10:11-0400
"\eta=\frac{0.5mv^2}{mgh}\\\eta=\frac{0.5(10)^2}{(9.8)(9)}=0.57"